# Enhancing Fairness in Unsupervised Graph Anomaly Detection through Disentanglement

### v1

@@ 3.3. Graph Anomaly Detection
In this phase, the primary objective is to detect anomalies unbiased to any demographic groups based on deterministic sensitive-irrelevant representations $\bar{\mathbf{Z}}_{x}=\bm{\mu}$. Due to anomalies typically significantly deviate the majority of samples, the reconstruction error is widely used to measure anomaly scores in prior studies (Ding et al., 2019; Fan et al., 2020; Shekhar et al., 2021). Thus, we introduce an additional decoder $f_{\phi}$ to evaluate anomalous node by reconstructing the input attributes $\mathbf{X}$ from sensitive-irrelevant representations $\bar{\mathbf{Z}}_{x}$. The anomaly score ${o}_{i}$ for node $v_{i}$ is defined as the corresponding reconstruction error:
(17) ${o}_{i}={||\mathbf{x}_{i}-\tilde{\mathbf{x}}_{i}||}_{F}^{2},$
where $\tilde{\mathbf{X}}=f_{\phi}(\bar{\mathbf{Z}}_{x},\tilde{\mathbf{Z}}_{s})$ and $\tilde{\mathbf{Z}}_{s}$ denotes a shuffled variant of $\mathbf{Z}_{s}$. Considering the potential bias in graph topology, we solely reconstruct node attributes to evaluate anomalies and employ MLP as the backbone of $f_{\phi}$ to mitigate the impact of biased topology during the message-passing process.
@@
Additionally, given that input attributes are likely to be relevant to sensitive attributes (Deepak and Abraham, 2020) while $\bar{\mathbf{Z}}_{x}$ is designed to be devoid of sensitive information, the reconstruction error inherently correlates with sensitive attributes. To mitigate the impact of this correlation and prevent directly leveraging sensitive attributes, we propose a correlation constraint term, which measures the absolute correlation between the reconstruction error ${o}_{i}$ and predicted sensitive attributes $\mathbf{z}_{s}^{i}$ as follows:
(18) $\mathcal{L}_{corr}=\left|\frac{(\sum_{i\in\mathcal{V}}{{o}_{i}-\mu_{o}})(\sum_{i\in\mathcal{V}}{\mathbf{z}_{s}^{i}-\mu_{zs}}))}{\sigma_{o}\sigma_{zs}}\right|,$
@@
where $\mu_{o}$ and $\mu_{zs}$ represent the means, while $\sigma_{o}$ and $\sigma_{zs}$ denote the corresponding standard deviations of $\mathbf{o}$ and $\mathbf{Z}_{s}$, respectively.
3.3.1. Learning Objective.
@@
In the anomaly detection phase, the encoder $f_{e}$ is set to be non-trainable, as it has already mastered the separation of sensitive-relevant and sensitive-irrelevant representations during the disentangled representation learning phase. Accordingly, the optimization is exclusively concentrated on the decoder $f_{\phi}$. The overall learning objective in the anomaly detection phase, consisting of the reconstruction term and the correlation constraints term can be formulated as:
(19) $\mathcal{L}_{ad}=\mathcal{L}_{rec}^{X}+\beta\mathcal{L}_{corr},$

### v2

@@ 3.2 Graph Anomaly Detection
Next, the primary objective is to detect anomalies unbiased to any demographic groups based on deterministic sensitive-irrelevant representations $\bar{\mathbf{Z}}_{x}=\bm{\mu}$. Since anomalies typically significantly deviate from the majority of nodes, reconstruction error has been widely used to measure anomaly scores (Ding et al., 2019; Fan et al., 2020; Shekhar et al., 2021). Considering the potential bias in graph topology as demonstrated in previous studies (Rahman et al., 2019; Spinelli et al., 2021; Zhu et al., 2023), we solely reconstruct node attributes and employ MLP as the backbone of the attribute decoder $f_{\phi}$ to mitigate the impact of biased topology during the message-passing process. The anomaly score ${o}_{i}$ is evaluated by reconstructing node attributes $\mathbf{X}$ from sensitive-irrelevant representations $\bar{\mathbf{Z}}_{x}$:
${o}_{i}={||\mathbf{x}_{i}-\tilde{\mathbf{x}}_{i}||}_{F}^{2},$ (17)
@@
where $\tilde{\mathbf{X}}=f_{\phi}(\bar{\mathbf{Z}}_{x},\tilde{\mathbf{Z}}_{s})$ and $\tilde{\mathbf{Z}}_{s}$ denotes a shuffled variant of $\mathbf{Z}_{s}$. Additionally, since sensitive attributes are correlated with observed node attributes (Deepak & Abraham, 2020; Sarhan et al., 2020; Oh et al., 2022) and $\bar{\mathbf{Z}}_{x}$ is supposed to be devoid of sensitive information, the reconstruction error inevitably correlates with sensitive attributes. To mitigate the impact of this correlation and prevent directly leveraging sensitive attributes, we propose a correlation constraint term, which measures the absolute correlation between the reconstruction error ${o}_{i}$ and predicted sensitive attributes $\mathbf{z}_{s}^{i}$:
$\mathcal{L}_{corr}=\left|\frac{(\sum_{i\in\mathcal{V}}{{o}_{i}-\mu_{o}})(\sum_{i\in\mathcal{V}}{\mathbf{z}_{s}^{i}-\mu_{zs}}))}{\sigma_{o}\sigma_{zs}}\right|,$ (18)
@@
where $\mu_{o}$ and $\mu_{zs}$ are the corresponding means, while $\sigma_{o}$ and $\sigma_{zs}$ are the standard deviations of $\mathbf{o}$ and $\mathbf{Z}_{s}$.
3.2.1 Learning Objective
@@
In the anomaly detection phase, the encoder $f_{e}$ is set to be non-trainable, as it has already mastered the separation of sensitive-relevant and -irrelevant representations during the disentangled representation learning phase. Accordingly, the optimization is exclusively concentrated on the decoder $f_{\phi}$. The overall loss consisting of the reconstruction term and the correlation constraints term can be formulated as:
$\mathcal{L}_{ad}=\mathcal{L}_{rec}^{X}+\beta\mathcal{L}_{corr},$ (19)